About 5% of the population has a particular genetic mutation. 400 people are randomly selected.
Find the mean for the number of people with the genetic mutation in such groups of 400.

Answer by ikleyn(52778)

.

    5% of 400 people is 0.05*400 = 5*4 = 20 people.    ANSWER

Solved.